Seafoam:

A company that manufactures storage bins for grains made a drawing of a silo. The silo has a conical base, as shown below: Which of the following could be used to calculate the total volume of grains that can be stored in the silo? π(2ft)2(8ft) + one over threeπ(2ft)2(9.5ft − 8ft) π(8ft)2(2ft) + one over threeπ(2ft)2(9.5ft − 8ft) π(2ft)2(8ft) + one over threeπ(9.5ft − 8ft)2(2ft) π(8ft)2(2ft) + one over threeπ(9.5ft − 8ft)2(2ft)

Seafoam:

1 attachment
Seafoam:

@AZ

AZ:

ah, that image was confusing for a sec but the bottom part is a cone So think of a shape like this

1 attachment
AZ:

We want to find the total volume so that means volume of cylinder + volume of cone do you know the formulas for volume of a cylinder or cone?

Seafoam:

V= pi 2 h?

Seafoam:

And for a cone its V= 1/ 3h pi r 2

AZ:

yeah For a cylinder V = \(\pi r^2 h\) For a cone V = \(\dfrac{1}{3} \pi r^2h\)

AZ:

so for the cylinder, we know the diameter is 4 what is the radius? and the height of the cylinder is 8 feet

Seafoam:

Do you have to calculate the cone and the cylinder seperately?

AZ:

yes, and we add them together to find the total volume

AZ:

Basically we break up the complex shape into shapes we know and calculate the volume and the total volume is the volume of each part all added together

Seafoam:

Right ok

AZ:

So you only have to plug in the numbers, you don't really need to calculate the final answers the radius of the cone is going to be the same radius as the cylinder diameter = radius * 2 we know the diameter is 4 so what is the radius?

AZ:

and the height of the cone is the difference of the total height and the cylinder's height we know the entire thing is 9.5 feet and the cylinder has a height of 8 so the height of the cone is what gets added to the cylinder's height to give you the total height of 9.5 so the height of the cone is 9.5 - 8 = ??

Seafoam:

Uh 1.5

Seafoam:

Wait how do you figure out the radius?

AZ:

The radius is half of the diameter 

Seafoam:

Oh 2 then

Seafoam:

Volumes 6.28 then rounded to 6.3

AZ:

exactly so total volume = volume of cylinder + volume of cone total volume = \(\pi r^2h_1 + \dfrac{1}{3} \pi r^2 h_2\) plug in r = 2 the first height is 8 the second height is 1.5

AZ:

Look at your answer choices, we don't need to actually calculate the volume we just need to set it up and I realized that we don't need to do the subtract either the height of the cone should have just been left as (9.5 - 8)

Seafoam:

Oh so is it the first option?

AZ:

Yup!

Seafoam:

Alright great